OpenStudy (anonymous):

(9x+2)/(x-9) = (2/5)

hero (hero):

Cross multiply

hero (hero):

\[\frac{9x + 2}{x - 9} = \frac{2}{5}\] 5(9x + 2) = 2(x - 9) Distribute: 45x + 10 = 2x - 18 Finish solving from there.